OpenStudy (anonymous):

What is the decay rate of the function y=0.5(0.83)^x

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

@Mertsj @abb0t @amistre64

OpenStudy (anonymous):

maybe they can help you :)

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

lol thanks! my final is tomorrow so i'm freaking out. thanks for trying though! :)

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

well there's all my smart friends :) goodluck and your welcome

OpenStudy (amistre64):

the thing being exponentiated i believe

OpenStudy (amistre64):

think of it like an interest rate ....

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

Like over a period of time?

OpenStudy (amistre64):

yes, "x" refers to time span

OpenStudy (anonymous):

so you make an x and y graph type thing and sort of resolve for each one?

OpenStudy (amistre64):

p0 = k(r)^0 p1 = k(r)^1 p2 = k(r)^2 the rate is the value being ^-ed

OpenStudy (anonymous):

ok.

OpenStudy (amistre64):

if |r| > 1 it grows if |r| < 1 it decays

OpenStudy (anonymous):

but then don't you have to find at what rate it decays at?

OpenStudy (amistre64):

it already gives you its rate ....

OpenStudy (anonymous):

on my answer sheet the teacher gaves us it says the answer isr=0.17 i just dont know how she got that.

OpenStudy (amistre64):

y = 0.5 (0.83)^x starting amount rate

OpenStudy (anonymous):

*is r=0.17

OpenStudy (amistre64):

let me see if im misreading this .....

OpenStudy (anonymous):

ok! :)

OpenStudy (anonymous):

thank you by the way.

OpenStudy (amistre64):

i think i see it now the rate of change is expressed as: a (1+r)^x 1+r = .83 r = .83 - 1 r = -.17 the negative just means its decaying instead of growing

OpenStudy (anonymous):

ok. thank you!!! :)

OpenStudy (amistre64):

good luck :)